OpenStudy (anonymous):

please could someone help to find the closed form of sequence 2.3, 1.6, 0.9, 0.2

OpenStudy (cwrw238):

i think you mean what is the formula for the nth term right?

OpenStudy (cwrw238):

a(n) =

OpenStudy (cwrw238):

a1 = 2.3 n=1 a2 = 1.6 n=2 a3 = 0.9 n =3 a4 = 0.2 n = 4 we need to relate n to these numbers subtraction is involved a1 = 3.0 - 7(0.1) a2 = 3.0 - 7(0.2) etc but we need to involve n = 1, 2 etc a(n) = 0.1[(30 - 7(n)] - this is the closed form

OpenStudy (anonymous):

thank you that make sense now

OpenStudy (cwrw238):

your welcome